john's age and mary's age sum up to 48. 60years later,john's age was double that of mary.
what are thier ages?

Found 2 solutions by fractalier, macston:
Answer by fractalier(6550)

Call John and Mary's age NOW, j and m. Then we have
j + m = 48, or j = 48 - m
Sixty years later,
j + 60 = 2(m + 60)
Now substitute in for j and get
48 - m + 60 = 2(m + 60)
108 - m = 2m + 120
This doesn't work...I think you meant six years later...
j + 6 = 2(m + 6)
48 - m + 6 = 2(m + 6)
54 - m = 2m + 12
42 = 3m
m = 14
j = 34

Answer by macston(5194)

.
J=John's age; M=Mary's age
.
J+M=48
J=48-M
.
J+60=2(M+60)
48-M+60=2M+120
108-M=2M+120
-12=3M
-4=M
ANSWER 1: Mary is -4 years old.
.
J=48-M=48-(-4)=52
ANSWER 2: John is 52 years old.
.
CHECK:
J+60=2(M+60)
52+60=2(-4+60)
112=2(56)
112=112
.
There is no reasonable answer for this problem.
Please check your problem.